OpenStudy (anonymous):

Describe how to transform the graph of f into the graph of g. f(x) = square root of quantity x minus nine. and g(x) = square root of quantity x plus five.

OpenStudy (anonymous):

@iGreen any help on this one?

OpenStudy (igreen):

\(f(x) = \sqrt{x-9}\) \(g(x) = \sqrt{x+5}\) -9 + ? = 5

OpenStudy (igreen):

@sunshine201

OpenStudy (anonymous):

14

OpenStudy (anonymous):

But is that to the right or left?

OpenStudy (igreen):

Right so it's 14. If you added 14 to 9 in the f(x), f(x) will become g(x).

OpenStudy (igreen):

What do you mean?

OpenStudy (anonymous):

The questions says to describe the transform. I get that it's 14 units but the answers ask for to the right or left. I'll post the answer choices.

OpenStudy (anonymous):

Shift the graph of f right 14 units. Shift the graph of f right 4 units. Shift the graph of f left 14 units. Shift the graph of f left 4 units.

OpenStudy (igreen):

The graph f(x) is to the right of g(x) by 14. So what direction will you shift it?

OpenStudy (anonymous):

left?

OpenStudy (igreen):

Yep! So your answer's gonna be shift 14 units left.

OpenStudy (anonymous):

Thank you so much!!